To break even in a manufacturing​ business, income or revenue R must equal the cost of production C. The cost C to produce x skateboards is C=63+22x. The skateboards are sold wholesale for ​$25 ​each, so revenue R is given by R=25x. Find how many skateboards the manufacturer needs to produce and sell to break even.​ (Hint: Set the cost expression equal to the revenue expression and solve for​ x.)


Question content area bottom


Part 1


The manufacturer needs to produce and sell

enter your response here


skateboards to break even.

​(Simplify your​ answer.)


1
Expert's answer
2022-01-18T17:50:54-0500

C=63+22x

R=25x

To break even, C=R

63+22x=25x

Group like terms,

25x-22x=63

3x=63

Simplify to get,

x=21

Thus to break even the manufacturer needs to produce 21 skateboards.